The Head of the Mathematics Department is interested to know whether there is any variation in the way the two instructors in the deprtment grade mathematics quiz for the mid-term examination. They were given the same set of 30 exam papers selected randomaly with the same marking scheame. The first instructor's grades have a variance of 52.3. The second instructor's grades have a variance of 89.9. Test the claim that the first instructor's variance is smaller at the 10% level of significance.
1
Expert's answer
2020-06-16T18:03:30-0400

The provided sample variances are "s_1^2=52.3" and "s_2^2=89.9" and the sample sizes are given by "n_1=30" and "n_2=30."

The following null and alternative hypotheses need to be tested:

"H_0:\sigma_1^2\geq \sigma_2^2"

"H_1:\sigma_1^2< \sigma_2^2"

This corresponds to a left-tailed test, for which a F-test for two population variances needs to be used.

Based on the information provided, the significance level is "\alpha=0.1," and the the rejection region for this left-tailed test is "R=\{F:F_L<0.617\}."

The F-statistic is computed as follows:


"F=\dfrac{s_1^2}{s_2^2}=\dfrac{52.2}{89.9}\approx0.582"

Since from the sample information we get that  "F=0.582<0.617=F_L," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the first instructor's variance of 52.3 is less than the second instructor's variance​ of 89.9, at the 10% level of significance.